Allison:

Autumn decides she’s going to count every leaf that she rakes from her yard. Unfortunately she forgets to count as she rakes. She knows that she started with 1000 leaves, and she counts 135 leaves that when she decides she is done. What is the percent change of number of leaves that Autumn raked?

Hero:

\(\%\triangle = \dfrac{a - b}{a} \times 100\) \(a = \text{original amount}\) \(b = \text{other amount}\)

Allison:

Yeah I got that

Allison:

Oh wait

Allison:

1000-135/1000?

Hero:

You forgot the times 100 part

Allison:

1000-135/1000*100

Allison:

There?

Hero:

You can't write it like that because it makes it look like you're writing \(\dfrac{1000 - 135}{1000*100}\)

Allison:

I can’t do Latex oml

Hero:

Shall I link you to LaTeX practicing?

dude:

She's on mobile, for the meantime you can do parenthesis instead (1000-135/1000)*100

Hero:

Still incorrect. That could be interpreted as \(\left(1000 - \dfrac{135}{1000}\right)*100\)

Allison:

No Hero, I do HTML here

Hero:

You have to learn how to either do \(\LaTeX\) or type equations linearly.

Elsa213:

\(\LARGE\frac{1000-135}{1000*100}\)

Hero:

See even @Elsa213 can do it

Hero:

Even though that's wrong too.

Elsa213:

I tri. e.e

Elsa213:

\(\bf\frac{1000-135}{1000}*100\) Der?

Allison:

I know what you were saying, I just typed it in the way you say is wrong. >.>

Hero:

\(\LaTeX\): \(\dfrac{1000 - 135}{1000} \times 100\) Linear method of the same expression: ((1000 - 135)/(1000)) x 100

Allison:

I got 1.156

Hero:

Show how you got that. I get something vastly different.

Allison:

1000-135/1000 is 1.156......

Hero:

If you say so. I greatly disagree though if you type what I've written above which is different from what you're typing.

Allison:

What ._.

dude:

You calculated incorrectly is what he means

dude:

You are supposed to subtract 1000 and 135 first

Allison:

But I didn’t

Allison:

865

dude:

divide that by 1000

Allison:

1.156

dude:

Dividing by any power of 10 only moves the decimal o.e Not sure what you're using x6zy3GZaRaC2mgQgFYIJvQ.png

Allison:

Oh I put 1000/856 o.e

dude:

Yeah the 1000-135 was at the numerator x.x

Allison:

e.e

Allison:

Stupid math, I swear.

dude:

You're level `99` ;p

Allison:

Yeh and?

Allison:

Cause it’s a number? o.e

dude:

You just haven't found your way of understanding math, like math other people :P

Allison:

Rip e.e

dude:

many instead of second math

Allison:

Jay

Allison:

Is it 856

Hero:

Let's start over and do it this way: First perform the following operation 1000 - 135 Let me know what you get.

Allison:

865

Hero:

Okay, take that amount and divide it by 1000. Let me know what you get.

Allison:

0.856

Hero:

Take that amount and multiply it by 100. Let me know what you get.

Allison:

By a thousand or a hundred?

dude:

Allison check again, it is not 0.856

Hero:

Does 100 look like hundred or thousand to you?

Allison:

e.e

dude:

its 0.865 you mixed the 6 and 5

Hero:

Anyway, at this point, I think you're toying around.

Hero:

I'm out

Allison:

K.

dude:

Multiply 0.865 by 100 C0t1fgSxQYKFuWU8ecPYyw.png

Allison:

Alright.

dude:

And thats your answer